I have a question in complex numbers in algebra that is what is the value of (w4+w5+w6)=?

Grade:9

1 Answers

ARc
6 years ago
As we know that (1+w+w^2)=0,w^4 + w^5 + w^6=w^4 (1+w+w^2)=w^4 x 0=0Here w and w^2 is the complex root of unitt